II- Translating verbal sentences into mathematical expression
1. The sum of two integers is 76 and their difference is 2. What are the integers?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)

The sentence is not ambiguous, and is precise. Just follow exactly what is written.

Answer by ikleyn(52776)

.
Let x and y be the numbers, x is the greater, y is the smaller.


Then the seeking translation is


    x + y = 76

    x - y =  2.


To solve this system of equations, add them. You will get


    2x = 76 + 2 = 78,   x = 78/2 = 39.


So, the greater number is 39.


Then you easily can find the smaller number: it is  2 units less, i.e.  y = 39-2 = 37.


ANSWER.  The numbers are  39  and 37.

Solved and thoroughly explained.